bemine11:

An isosceles triangle has one angle with a measure of 30°. What could the measure of the other two angles be? 30° and 30° 30° and 120° 75° and 75° 30° and 100° *I think it is 75 to 75*

bemine11:

IT is more then one

jhonyy9:



jhonyy9:

@Laylalyssa

jhonyy9:

questions ?

bemine11:

I dont get how to find the x

jhonyy9:

hope you know that an isoscel triangle has two sides and the corresponding two angles congruent yes ?

bemine11:

Yea i know that

jhonyy9:

and hope you know that the sum of angle measures of a triangle is equal 180 degrees yes ?

bemine11:

yea

bemine11:

I know everything about then i think

jhonyy9:

ok given that one angle measure of this triangle from your posted problem has 30 degrees so 180-30=150 what mean that the sum of other two angles of the presently triangle is 150 degrees but we know that this 150 degrees is the sum of two congruent triangles so x +x = 150 how you get the x value ?

jhonyy9:

@supie

jhonyy9:

x +x = 2x 2x = 150 divide both sides by 2 and get 2x/2 = 150/2 x = 75 do you understand it now ?

bemine11:

Yea

jhonyy9:

ok hnd cya

jhonyy9:

@Laylalyssa

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 @Laylalyssa \(\color{#0cbb34}{\text{End of Quote}}\) yes

jhonyy9:

do you know this way to solve the posted problem ?